OpenStudy (anonymous):

-1/2 * 3/8? I need to find the sum and simplify if possible so can someone PLEASE walk me through the correct steps?

OpenStudy (anonymous):

Cross multiply

OpenStudy (anonymous):

-1 x 3 ------ 2 x 8

OpenStudy (anonymous):

\(\Huge\frac{-1\ x\ 3}{2\ x\ 8}\)

OpenStudy (anonymous):

Alright and this is right because I don't think I can afford to mess up

OpenStudy (anonymous):

So what did you get?

OpenStudy (anonymous):

\(\Huge\frac{-1\times 3}{2 \times 8}\)